# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ Abstract
Methods. We simulate the evolution of irradiated giant planets with MESA, spanning the observed range of masses and equilibrium temperatures, incorporating an internal source of Ohmic dissipation that extends to deep layers of the envelope. We consider the heat-flux-dependent evolution of the deep-seated dynamo field, on which the induced field depends; we adopt a state-of-the-art electrical conductivity, considering the thermal ionization of alkali metals in the outer layers and the pressure-ionization in the interior, and the corresponding solutions for the induced currents across the planet.
Results. We infer average atmospheric wind intensities, averaged in the region $p\lesssim 10$ bar, in the range 0.01-1 km/s in order to reproduce the range of observed radii, decreasing roughly linearly with planetary mass, and much more steeply with equilibrium temperature. This is consistent with the expected effects of magnetic drag from the induced field, which is higher for more intense irradiation, via conductivity, and for larger masses, which have higher dynamo fields. Due to the evolution of the dynamo field and the proportionality of the induced currents on it, the Ohmic efficiency typically decreases by at least one order of magnitude from 0.1 to 10 Gyr, at contrast with the common assumption of a constant-in-time value. Notably, the extent of the main convective region, and the associated heat flux supporting the dynamo, is reduced in the presence of strong Ohmic dissipation, which in turn depends on the dynamo field strength, generating a non-trivial coupling of the latter with the atmospheric induction, potentially leading to an oscillatory behaviour of the field strength. These findings remain generally valid even when accounting for a long-term increase in the main-sequence host star luminosity, although this case can more readily lead to HJ re-inflation, consistent with previous studies.
@@ 1 Introduction
The intriguing exoplanet class of hot Jupiters (HJs) currently consists of several hundred objects, discovered and characterized in terms of mass and radius. These gas giant planets have orbital periods of a few days, corresponding to separations within $\sim$0.1 AU from their host stars (for a review, see Heng & Showman 2015 and Fortney et al. 2021). Such a proximity implies tidally locking, with strong stellar irradiation of their day sides, which results in typical equilibrium temperatures of up to $3000$ K, with a few even hotter. The intense irradiation flux, coupled with the strong thermal gradients due to the tidal locking, gives rise to powerful atmospheric flows, with equatorial super-rotation and strong day-night flows which try to redistribute the heat across the planetary surface (e.g. Cho et al. 2008; Dobbs-Dixon & Lin 2008; Showman et al. 2009; Heng et al. 2011; Rauscher & Menou 2012; Perna et al. 2012; Rauscher & Menou 2013; Perez-Becker & Showman 2013; Parmentier et al. 2013; Rogers & Showman 2014; Showman et al. 2015; Kataria et al. 2015; Koll & Komacek 2018; Beltz et al. 2022; Komacek et al. 2022; Sengupta & Sengupta 2023).
@@
Among these proposed mechanisms, the latter has received considerable attention (Batygin & Stevenson, 2010; Batygin et al., 2011; Perna et al., 2010a, b; Huang & Cumming, 2012; Wu & Lithwick, 2013; Rauscher & Menou, 2013; Rogers & Showman, 2014; Rogers & Komacek, 2014; Ginzburg & Sari, 2016; Rogers & McElwaine, 2017; Hardy et al., 2022; Beltz et al., 2022; Benavides et al., 2022; Knierim et al., 2022). The deep-seated magnetic field, generated in the metallic hydrogen region (French et al., 2012), is a fundamental ingredient, since the atmospheric flow of thermally ionized particles (mostly free electrons from alkali metals) advects its field lines, twisting them considerably in case of high magnetic Reynolds numbers, as it arguably happens in a large fraction of HJs (Dietrich et al., 2022; Soriano-Guerrero et al., 2025). Although the shear and the corresponding magnetic field winding happen in the shallow atmospheric layers (at pressures $p\lesssim 10$ bar), the twisting of the lines propagates into deeper convective regions, as modelled in previous works for cold giant planets (Liu et al., 2008) and HJs (Perna et al., 2010a, b; Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013; Ginzburg & Sari, 2016), with different assumptions. This results in an additional heat source for the planet, which slows down the cooling and shrinking, leading to a larger radius than that expected from standard cooling. While the extent of the magnetic drag (Perna et al., 2010a) in reducing flow speeds (and hence, the magnitude of Ohmic dissipation) is still debated (Menou, 2012a; Knierim et al., 2022), this heating source appears consistent with observations (Laughlin et al., 2011; Thorngren & Fortney, 2018). There is also the possibility that more than one mechanism may operate simultaneously (Sarkis et al., 2021).
@@
The goal of this paper is to study the long-term evolution (on Gyr timescales) of HJs. Using a specifically modified version of the public code MESA [^1] https://github.com/MESAHub/mesa (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), we run a grid of simulations for a range of planetary masses, equilibrium temperatures, and intensity of the currents, varying also the composition and core masses. Similarly to e.g. Batygin et al. (2011); Wu & Lithwick (2013), we use a temperature- and density-dependent conductivity, which makes the total Ohmic dissipation rate vary in time.
@@
We introduce three main novelties. The first one is the proportionality of the induced currents with the deep-seated dynamo magnetic field, which evolves in time via widely-used scaling laws that account for the internal heat flux at a given age (Christensen et al., 2009). The second one is the parametrization of the complex, intrinsically 3D wind pattern in the uppermost layer, with the average wind velocity in the atmospheric region, with the aim of constraining its value. This approach is similar but complementary to what done by Knierim et al. (2022), who however used a constant dynamo magnetic field (like previous studies), and focused more on the exploration of the detailed wind profile in the atmospheric region.
@@
The third novelty is the inclusion of the most updated values of temperature- and density-dependent electrical conductivity, taking into account both the thermal ionization of alkali metals, dominant at $\rho\lesssim 10^{-2}$  g cm-3(Kumar et al., 2021), and the latest calculation of electron conductivity arising from the pressure-ionization of H and He (Bonitz et al., 2024), important at higher densities. Finally, we propose the use of semi-analytical piecewise power law solutions for the current density profiles, applicable to arbitrary combinations of magnetic and wind velocities.
@@
The paper is organized as follows: Sect. 2 briefly presents the observational data; Sect. 3 describes the standard evolutionary models for HJs as implemented in MESA; Sect. 4 presents our Ohmic heating model and its major ingredients; Sect. 5 shows representative results of the simulations and discusses some implications; a discussion with the main conclusions is provided in Sect. 6.
@@
Figure 1: HJ inflated radii observational trends, for the 424 Jupiter-like exoplanets in the database with available estimated mass $M\geq 0.5\leavevmode\nobreak\ M_{j}$, age $t\geq 100$ Myr, radius $R$ and equilibrium temperature $T_{\rm eq}$. We consider only the planets with relative errors in radius of less than $25\%$ and available uncertainties for the mass. Here we have also excluded seven planets younger than 100 Myr. Top: planetary radius $R$ versus $T_{\rm eq}$ or, equivalently, irradiated flux $F_{\rm irr}=4\sigma_{\rm sb}T_{\rm eq}^{4}$. The colours of the data points identify the mass by ranges as follows: purple [0.5,0.8) $M_{j}$, dark blue [0.8,1.5) $M_{j}$, blue [1.5,2.5) $M_{j}$, green [2.5,4) $M_{j}$, yellow [4,8) $M_{j}$ and orange [8,13] $M_{j}$. With the same colour coding, the lines represent the best-fit reported in Table 2, for the planets within each mass range, and considering only $T_{\rm eq}>1000$ K. For the sake of better visibility of the bulk of planets, we left outside the plot three ultra-HJ with $T_{\rm eq}>3300$ K, but they are considered in the statistical trends. The grey background boxes indicate the mean and standard deviation of the observed radius $\bar{R}\pm\delta_{R}$ for all HJs within a 250 K-wide ranges of $T_{\rm eq}$, for a range of $T_{\rm eq}$ for which the inflation starts to appear, and with enough planets per bin. Bottom: Histograms of the observed radii for the same $T_{\rm eq}$ ranges, per each $T_{\rm eq}$ range of the grey boxes of the top panel. Due to the small size of the hottest HJ sample, we group two $T_{\rm eq}$ bins in the last histogram.
@@ 2 Observational sample of Hot Jupiters
Here we present a short overview of the observational data on HJs, obtained from the NASA Exoplanet Archive. As of June 1st, 2025, there were a total of 1899 confirmed gas giants with known masses in the range considered here, $M\in[0.5,13]\leavevmode\nobreak\ M_{j}$ (we denote the Jovian mass and radius by $M_{j}$ and $R_{j}$, respectively). The upper limit of the range is simply the Deuterium-burning boundary with brown dwarfs. The lower limit is conservatively in agreement with other studies, e.g. Sestovic et al. (2018), who excluded the lightest planets whose masses and radii might be affected by mass loss.
Table 1: Observed HJs trends as a function of mass range.
Mass [$M_{j}$] Number $A$ $C$
@@
 [^2] Columns indicate: number of objects with $T_{\rm eq}\geq 1000$ K and best-fit parameters of the functional form $R/R_{j}=C+A\log(F/F_{s})$, where $F_{s}$ corresponds to the irradiation flux of $T_{\rm eq}=1000$ K.
@@
Since the radius is our main observable to compare with, we have excluded data with high errors in radius, and only selected those planets with errors less than $25\%$. Moreover, we have discarded those planets with only upper limits on the mass and without associated mass uncertainties. We also exclude 7 planets younger than $100$ Myr, since at those ages the planet shrinking has some memory of the initial entropy and the very early ages, which is unknown, therefore introducing additional free parameters in the problem.
@@
These selection criteria yield a data set of 424 gas giants. Fig. 1 shows the measured values of $R$ versus $T_{\rm eq}$ (top panel) and the distribution of radii for different planets grouped by different ranges of temperature with 250 K-wide bins centred around values in the range $T_{\rm eq}=[1000,2500]$ K (six small panels in the bottom), with the associated mean and standard deviation values, $\bar{R}$ and $\sigma_{R}$ represented by grey boxes in the top panel. As is well known (e.g., Menou 2012a), the radii of the planets increase as the equilibrium temperature rises, for planets with $T_{\rm eq}\gtrsim 1000$ K. The average radius increases up to $\bar{R}=1.48\pm 0.18\leavevmode\nobreak\ R_{j}$ for the $T_{\rm eq}\in[1875,2125]$ K sample, and up to $\bar{R}=1.53\pm 0.19\leavevmode\nobreak\ R_{j}$ for the $T_{\rm eq}\in[2125,2375]$ K sample (which is, however, sparsely populated). The coloured lines in the figure show the best-fitting trends of the form $R/R_{j}=C+A\log(F_{\rm irr}/F_{s})$, where we fix $F_{s}=0.227$ MW m-2, i.e. the irradiation flux corresponding to $T_{\mathrm{eq}}=1000$ K. The fit is performed for planets with $T_{\mathrm{eq}}\gtrsim 1000$ K, across different mass ranges, as reported in Table 2. Such fits are comparable to the values provided by Sestovic et al. (2018), who fitted the same functional form for planets with $F\geq F_{s}$, but they let $F_{s}$ to vary. Other older studies (Laughlin et al., 2011; Enoch et al., 2012) use different functional forms, all non-linear with $T_{\rm eq}$ as well, with a sample a few times smaller. Considering the differences in samples and functional forms, our results can be considered overall consistent.
@@
These trends show that lighter HJs ($M\sim 0.5-1.5\leavevmode\nobreak\ M_{j}$) have a statistically significant steeper dependence with irradiation ($A\sim 0.5$), compared to the heavier ones. The extremely massive HJs, $M\gtrsim 8\leavevmode\nobreak\ M_{j}$ (close to the border with brown dwarfs) barely show signs of inflation. Such trend with mass is well known, (e.g. Sestovic et al. 2018; Thorngren 2024 and references within), and, at first level, can be understood by the gravity counteracting inflation, for a given rate of deposited heat (see also e.g. Lopez & Fortney 2016 for a brief discussion).
@@
Below we assess this dispersion in terms of the most important factors intervening in the planetary long-term cooling, considering the dependence of the modelled radius evolution on observables ($T_{\rm eq},M$) and on the deposition of heat. Factors like composition, mass core can also contribute to the dispersion of the sample, but in this study we will not explore them. We refer the reader to App. C for a brief exploration of composition, core mass and density, with references to the extended literature on this aspect.

### v2

@@ Abstract
Methods. Using MESA, we simulated the evolution of irradiated giant planets spanning the observed range of masses and equilibrium temperatures, incorporating an internal source of Ohmic dissipation that extends to deep layers of the envelope. We considered the heat-flux-dependent evolution of the deep-seated dynamo field on which the induced field depends. We adopted a state-of-the-art electrical conductivity, considering the thermal ionisation of alkali metals in the outer layers and the pressure-ionisation in the interior and the corresponding solutions for the induced currents across the planet.
Results. We inferred that, in order to reproduce the range of observed radii, the atmospheric wind intensities averaged in the region $p\lesssim 10$ bar have to be in the range 0.01-1 km/s and to decrease roughly linearly with planetary mass and much more steeply with equilibrium temperature. This is consistent with the expected effects of magnetic drag from the induced field, which is higher for more intense irradiation, via conductivity, and for larger masses, which have higher dynamo fields. Due to the evolution of the dynamo field and the proportionality of the induced currents on it, the Ohmic efficiency typically decreases by at least one order of magnitude from 0.1 to 10 Gyr, which is in contrast with the common assumption of a constant-in-time value. Notably, the extent of the main convective region and the associated heat flux supporting the dynamo is reduced in the presence of strong Ohmic dissipation, which in turn depends on the dynamo field strength, generating a non-trivial coupling of the latter with the atmospheric induction and potentially leading to the oscillatory behaviour of the field strength. These findings remain generally valid even when accounting for a long-term increase in the main-sequence host star luminosity, although this case can more readily lead to HJ re-inflation, consistent with previous studies.
@@ 1 Introduction
The intriguing exoplanet class of hot Jupiters (HJs) currently consists of several hundred objects, discovered and characterised in terms of mass and radius. These gas giant planets have orbital periods of a few days, corresponding to separations within $\sim$0.1 AU from their host stars (for a review, see Heng & Showman 2015 and Fortney et al. 2021). Such proximity implies tidal locking, with strong stellar irradiation of their day sides, which results in typical equilibrium temperatures of up to $3000$ K, with a few that are even hotter. The intense irradiation flux coupled with the strong thermal gradients due to the tidal locking gives rise to powerful atmospheric flows, with equatorial super-rotation and strong day-night flows that try to redistribute the heat across the planetary surface (e.g. Cho et al. 2008; Dobbs-Dixon & Lin 2008; Showman et al. 2009; Heng et al. 2011; Rauscher & Menou 2012; Perna et al. 2012; Rauscher & Menou 2013; Perez-Becker & Showman 2013; Parmentier et al. 2013; Rogers & Showman 2014; Showman et al. 2015; Kataria et al. 2015; Koll & Komacek 2018; Beltz et al. 2022; Komacek et al. 2022; Sengupta & Sengupta 2023).
@@
Among these proposed mechanisms, Ohmic dissipation has received considerable attention (Batygin & Stevenson, 2010; Batygin et al., 2011; Perna et al., 2010a, b; Huang & Cumming, 2012; Wu & Lithwick, 2013; Rauscher & Menou, 2013; Rogers & Showman, 2014; Rogers & Komacek, 2014; Ginzburg & Sari, 2016; Rogers & McElwaine, 2017; Hardy et al., 2022; Beltz et al., 2022; Benavides et al., 2022; Knierim et al., 2022). The deep-seated magnetic field, generated in the metallic hydrogen region (French et al., 2012), is a fundamental ingredient since the atmospheric flow of thermally ionised particles (mostly free electrons from alkali metals) advects its field lines, twisting them considerably in case of high magnetic Reynolds numbers, as it arguably happens in a large fraction of HJs (Dietrich et al., 2022; Soriano-Guerrero et al., 2025). Although the shear and the corresponding magnetic field winding happen in the shallow atmospheric layers (at pressures $p\lesssim 10$ bar), the twisting of the lines propagates into deeper convective regions, as modelled in previous works for cold giant planets (Liu et al., 2008) and HJs (Perna et al., 2010a, b; Batygin & Stevenson, 2010; Batygin et al., 2011; Wu & Lithwick, 2013; Ginzburg & Sari, 2016), with different assumptions. This results in an additional heat source for the planet that slows down the cooling and shrinking, leading to a larger radius than that expected from standard cooling. While the extent of the magnetic drag (Perna et al., 2010a) in reducing flow speeds (and hence the magnitude of Ohmic dissipation) is still debated (Menou, 2012a; Knierim et al., 2022), this heating source appears consistent with observations (Laughlin et al., 2011; Thorngren & Fortney, 2018). There is also the possibility that more than one mechanism may operate simultaneously (Sarkis et al., 2021).
@@
The goal of this paper is to study the long-term evolution (on gigayear timescales) of HJs. Using a specifically modified version of the public code MESA [^1] https://github.com/MESAHub/mesa (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023), we ran a grid of simulations for a range of planetary masses, equilibrium temperatures, and intensity of the currents, varying also the composition and core masses. Similarly to Batygin et al. (2011); Wu & Lithwick (2013), for example, we used a temperature- and density-dependent conductivity, which makes the total Ohmic dissipation rate vary in time.
@@
We introduce three main novelties. The first is the proportionality of the induced currents with the deep-seated dynamo magnetic field, which evolves in time via widely used scaling laws that account for the internal heat flux at a given age (Christensen et al., 2009). The second is the parametrisation of the complex, intrinsically 3D wind pattern in the uppermost layer, with the average wind velocity in the atmospheric region. Our aim is to constrain its value. This approach is similar but also complementary to what has been done by Knierim et al. (2022), who nevertheless used a constant dynamo magnetic field (like previous studies) and focused more on the exploration of the detailed wind profile in the atmospheric region.
@@
The third novelty is the inclusion of the most updated values of temperature- and density-dependent electrical conductivity, taking into account both the thermal ionisation of alkali metals, dominant at $\rho\lesssim 10^{-2}$  g cm-3(Kumar et al., 2021), and the latest calculation of electron conductivity arising from the pressure-ionisation of H and He (Bonitz et al., 2024), which are important at higher densities. Finally, we propose the use of semi-analytical piecewise power-law solutions for the current density profiles, applicable to arbitrary combinations of magnetic and wind velocities.
@@
The paper is organised as follows: Sect. 2 briefly presents the observational data. In Sect. 3 we describe the standard evolutionary models for HJs as implemented in MESA. Sect. 4 presents our Ohmic heating model and its major ingredients. Sect. 5 shows representative results of the simulations, and we discusses some implications. We provide a discussion and the main conclusions in Sect. 6.
@@
Figure 1: Hot Jupiter inflated radii observational trends for the 424 Jupiter-like exoplanets in the database with available estimated mass $M\geq 0.5\leavevmode\nobreak\ M_{\rm j}$; age $t\geq 100$ Myr; radius, $R$; and equilibrium temperature, $T_{\rm eq}$. We consider only the planets with relative errors in a radius of less than $25\%$ and available uncertainties for the mass. We also excluded seven planets younger than 100 Myr. Top: Planetary radius, $R$, versus, $T_{\rm eq}$, or, equivalently, irradiated flux $F_{\rm irr}=4\sigma_{\rm sb}T_{\rm eq}^{4}$. The colour of the data points identify the mass by ranges as follows: purple [0.5,0.8) $M_{\rm j}$, dark blue [0.8,1.5) $M_{\rm j}$, blue [1.5,2.5) $M_{\rm j}$, green [2.5,4) $M_{\rm j}$, yellow [4,8) $M_{\rm j}$ and orange [8,13] $M_{\rm j}$. With the same colour coding, the lines represent the best-fit reported in Table 2, for the planets within each mass range, and considering only $T_{\rm eq}>1000$ K. For the sake of better visibility of the bulk of planets, we left out three ultra-HJ with $T_{\rm eq}>3300$ K from the plot, but they are considered in the statistical trends. The grey background boxes indicate the mean and standard deviation of the observed radius $\bar{R}\pm\delta_{R}$ for all HJs within a 250 K-wide range of $T_{\rm eq}$. Bottom: Histograms of the observed radii for the same $T_{\rm eq}$ ranges per each $T_{\rm eq}$ range of the grey boxes of the top panel. Due to the small size of the hottest HJ sample, we grouped two $T_{\rm eq}$ bins in the last histogram.
@@ 2 Observational sample of Hot Jupiters
Here we present a short overview of the observational data on HJs obtained from the NASA Exoplanet Archive. As of June 1, 2025, there were a total of 1899 confirmed gas giants with known masses in the range considered here, $M\in[0.5,13]\leavevmode\nobreak\ M_{\rm j}$ (we denote the Jovian mass and radius by $M_{\rm j}$ and $R_{\rm j}$, respectively). The upper limit of the range is simply the Deuterium-burning boundary with brown dwarfs. The lower limit is conservatively in agreement with other studies, such as Sestovic et al. (2018), who excluded the lightest planets whose masses and radii might be affected by mass loss.
Table 1: Observed HJ trends as a function of mass range.
Mass [$M_{\rm j}$] Number $A$ $C$
@@
 [^2] Columns indicate: mass ranges, number of objects with $T_{\rm eq}\geq 1000$ K and best-fit parameters of the functional form $R/R_{\rm j}=C+A\log(F/F_{s})$, where $F_{s}$ corresponds to the irradiation flux of $T_{\rm eq}=1000$ K.
@@
Since the radius is our main observable to compare with, we excluded data with high errors in radius and only selected those planets with errors of less than $25\%$. Moreover, we have discarded those planets with only upper limits on the mass and without associated mass uncertainties. We also exclude seven planets younger than $100$ Myr, since at those ages the planet shrinking has some memory of the initial entropy and the very early ages, which is unknown, therefore introducing additional free parameters in the problem.
@@
These selection criteria yield a data set of 424 gas giants. Fig. 1 shows the measured values of $R$ versus $T_{\rm eq}$ (top panel) and the distribution of radii for different planets grouped by different ranges of temperature with 250 K-wide bins centred around values in the range $T_{\rm eq}=[1000,2500]$ K (six small panels in the bottom), with the associated mean and standard deviation values, $\bar{R}$ and $\sigma_{R}$ represented by grey boxes in the top panel. As is well known (e.g. Menou 2012a), the radii of the planets increase as the equilibrium temperature rises, for planets with $T_{\rm eq}\gtrsim 1000$ K. The average radius increases up to $\bar{R}=1.48\pm 0.18\leavevmode\nobreak\ R_{\rm j}$ for the $T_{\rm eq}\in[1875,2125]$ K sample, and up to $\bar{R}=1.53\pm 0.19\leavevmode\nobreak\ R_{\rm j}$ for the $T_{\rm eq}\in[2125,2375]$ K sample (which is, however, sparsely populated). The coloured lines in the figure show the best-fitting trends of the form $R/R_{\rm j}=C+A\log(F_{\rm irr}/F_{s})$, where we fix $F_{s}=0.227$ MW m-2, i.e. the irradiation flux corresponding to $T_{\mathrm{eq}}=1000$ K. The fit is performed for planets with $T_{\mathrm{eq}}\gtrsim 1000$ K, across different mass ranges, as reported in Table 2. Such fits are comparable to the values provided by Sestovic et al. (2018), who fitted the same functional form for planets with $F\geq F_{s}$, but they let $F_{s}$ to vary. Other older studies (Laughlin et al., 2011; Enoch et al., 2012) use different functional forms, all non-linear with $T_{\rm eq}$ as well, with a sample a few times smaller. Considering the differences in samples and functional forms, our results can be considered overall consistent.
@@
These trends show that lighter HJs ($M\sim 0.5-1.5\leavevmode\nobreak\ M_{\rm j}$) have a statistically significant steeper dependence with irradiation ($A\sim 0.5$), compared to the heavier ones. The extremely massive HJs, $M\gtrsim 8\leavevmode\nobreak\ M_{\rm j}$ (close to the border with brown dwarfs) barely show signs of inflation. This trend with mass is well known, (e.g. Sestovic et al. 2018; Thorngren 2024 and references within) and can initially be understood by the gravity counteracting inflation for a given rate of deposited heat (see also e.g. Lopez & Fortney 2016 for a brief discussion).
@@
Below we assess this dispersion in terms of the most important factors intervening in the planetary long-term cooling, considering the dependence of the modelled radius evolution on observables ($T_{\rm eq},M$) and on the deposition of heat. Factors like composition, mass core can also contribute to the dispersion of the sample, but in this study we do not explore them. We refer the reader to App. C for a brief exploration of composition, core mass and density, with references to the extended literature on this aspect.